OpenStudy (anonymous):

What is the y-intercept of the line represented by the equation? y = 7x + 2 will fan and medal

OpenStudy (sleepyhead314):

slope intercept form is y = mx + b where b is the y-intercept y = mx + b y = 7x + 2 ^ take a guess? ;)

OpenStudy (anonymous):

y = mx + b Where 'b' is the y-intercept. So what's the y-intercept here?

OpenStudy (clongoria2):

The y-intercept is always going to be where the + or - is. So the y-intercept = ?

OpenStudy (anonymous):

Lagging comments..would've been the first :P >.> <.<

OpenStudy (anonymous):

oh duh its 2 XD

OpenStudy (anonymous):

You got it.

OpenStudy (clongoria2):

Not a chance :P

OpenStudy (clongoria2):

@iGreen.

OpenStudy (sleepyhead314):

good job ^_^ @bcimkenz

OpenStudy (anonymous):

Hehe, I type up to 100WPM, I think I do have a chance :P

OpenStudy (clongoria2):

@iGreen. Nice, i can only type 90wpm

OpenStudy (sleepyhead314):

I only type at max 45wpm and I still beat yall ;)

OpenStudy (anonymous):

'Cause you got here first, and my comments were lagging. When I hit the post button none of you typed a reply yet :P